OpenStudy (anonymous):

I need help solving this equation. 6r^2+11r=10. My answers are r+and- i5/2 squared and r+and- 2/3 squared...I am wrong apparently.

Directrix (directrix):

6r^2+11r=10 6 r^2 + 11r - 10 = 0 @Amandarangel18 Factor: 6 r^2 + 11r - 10 = 0. After that, you can use the Zero Product Property to get the roots. Please post what you get when you factor 6 r^2 + 11r - 10 = 0. Then, we can compare results.

Directrix (directrix):

@Amandarangel18 I need help solving this equation - I'm trying to help.

OpenStudy (anonymous):

On this problem you need to rearrange 6r^2+11r=10 to be 6r^2+11r-10=0 Then use the quadratic formula to solve. \[-b \pm \sqrt{b ^{2}-4ac}\div2a\] Which then gives: \[-11\pm \sqrt{11^{2}-4(6)(-10)}\div2(6)\] So r=2/3, -5/2 Plug those values back in to the original equation for r and you'll see the solution works out.

Directrix (directrix):

@jonnymiller OpenStudy values the Learning process - not the ‘Give you an answer’ process •Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

I did guide the asker to a solution!

OpenStudy (zzr0ck3r):

There is no way that anyone outside of @jonnymiller thinks that what he did was "to guide" ...

OpenStudy (anonymous):

In jonnymiller's defense, I already did the problem and finished it before I posted the equation. Hence, the answer in the question that I asked. We both got the same answer, but that answer is wrong apparently. According to my math book.

OpenStudy (anonymous):

I found out that I didnt rationalize which kept me from getting the correct answer. I missed the last step which was to rationalize.